OpenStudy (anonymous):

WHAT DOES X EQUAL 5/6 X 3/X=1/4

OpenStudy (anonymous):

Is it 5/6 * 3/x = 1/4

OpenStudy (anonymous):

whats the awnser 4 x

OpenStudy (anonymous):

If that is the case then 15/6x = 1/4 6x = 60 x = 10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2/3*9/x=3/11

OpenStudy (anonymous):

wat does x equal